Madden Company has projected its income before taxes for next year as shown below. Madden is subject to a 40% income tax rate.

Sales (160,000 units)

$ 8,000,000

Cost of sales

Variable costs

$2,000,000

Fixed costs

3,000,000

Total costs

(5,000,000)

Income before taxes

$ 3,000,000

Question

Madden’s net assets are $36,000,000. The dollar sales that must be achieved for Madden to earn a 10% after-tax return on assets would be

  • A.$8,800,000

  • B.$16,000,000

  • C.$12,000,000

  • D.$6,880,000
Answer #1
Madden company has projected its income before taxes
answered by: Jafer sherif Ahmed
Answer #2

Madden want to earn after tax return of 10% of net assets.

Return on net assets = Net Income/ Total assets

Total assets =$36000000

After tax return that Madden want is $36000000 * 10% = $3600000

Tax rate is 40%. So before tax income will be $3600000/60% = $6000000

Now Sales - Variable cost = Contribution margin.

Contribution margin - Fixed expenses = Income before Taxes

So Contribution margin for the required return = Income before taxes + Fixed expenses

It will be $6000000 + $3000000 = $9000000

Now Selling price per unit =$8000000/160000 units = $50 per unit

Varible expenses per unit = $2000000/160000 units = $12.5 per unit

Now let the number of units sold be x at the required sale amount.

So Contribution margin = 50x - 12.5x

Therefore $9000000 = 37.5x

So x = 240000 units

So Dollars sale will be 240000 units * $50 per unit

It will be $12000000

So our answer is option C i.e $12000000